Hoshi is working on an art project in the shape of a triangular prism. The sides of the triangle are 2.4 feet, 1.5 feet, and 1.3 feet. Find the measures of the angles of the triangle to the nearest tenth
Answer by addingup(3677)

Let's use the Law of Cosines to find one angle, and use it again to find a second angle. The third angle is easy because the angles in EVERY triangle add up to 180 degrees.
```````````````````````````````````````````````````````
Let's call the sides: 2.4 a, 1.5 b, and 1.3 c. And the angles A, B, and C:
cos(C) = (a^2+b^2−c^2)/2ab
cos(C) = (2.4^2+1.5^2-1.3^2)/[2(2.4*1.5)] = 0.8778
C = cos^-1(0.8778)= 28.62 degrees
```````````````````````````````````````````````````
Now you have angle C. Do the same for one other angle, then add the two angles and subtract the result from 180 to get your third angle.
John